OpenStudy (kenialove):

The product of three consecutive whole numbers is 990. What is the sum of the three whole numbers ? A. 30 B. 33 C. 63 D. 96

OpenStudy (mathstudent55):

Start with the sum of 30. Assume one number and start guessing.

OpenStudy (michele_laino):

hint: three consecutive natural numbers are like these: \[\large n,\quad n + 1,\quad n + 2\] so your problem reads now: \[\Large n \cdot \left( {n + 1} \right) \cdot \left( {n + 2} \right) = 990\]

OpenStudy (mathstudent55):

For example: Is 9 * 10 * 11 = 990? If not, Make a new assumption, and test it.

OpenStudy (mathstudent55):

For example: Is 8 * 9 * 10 = 990? If not, make a new assumption, and test it.